OpenStudy (anonymous):

@ganeshie8 The table shows the results of drawing one marble from a bag of colored marbles. Each marble was replaced after the outcome was recorded. If there are a total of 50 marbles in the bag, about how many of the marbles are green? Color Number of Pulls red 6 green 4 orange 7 black 3 A. 20 B. 8 C. 10 D. 4

OpenStudy (anonymous):

is it C

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@ganeshie8 am i right

ganeshie8 (ganeshie8):

Correct !

ganeshie8 (ganeshie8):

probability of green marble = 4/20 = 1/5

ganeshie8 (ganeshie8):

So expected value = 50*1/5 = 10